OpenStudy (anonymous):

If, x(t)=0.5(t)-4 where does x=0?

OpenStudy (anonymous):

t=8

OpenStudy (anonymous):

4/(1/2) =8

OpenStudy (anonymous):

makes sense

myininaya (myininaya):

if x=0, then we are looking for t such that it satisfies 0=.5t-4 add 4 on both sides 4=.5t 2(4)=t 8=t t=8 so when x=0, t=8

OpenStudy (anonymous):

hey myininaya, if you see this reply so I can ask you a more thorough question about this answer. Thanks!

OpenStudy (anonymous):

I am about to give up I am so frustrated, mainly because I feel I should know how to do this! The motion of a train is given by x(t)=.5t-3.5 create a position time graph with accurate slope and intercept. Done already. That was the easy part.  Find the position at 4 seconds. Confirm with an equation So, x(4)=0.5(4)-3.5 Solution: -1.5 At what instant is the position zero from the graph? I came up with 7 obviously. Then it says, from your equation, show your work. Now I am lost, help! lol

OpenStudy (anonymous):

for some reason it deleted half my post, what the heck

myininaya (myininaya):

so we want to know when x=0, what does t=? x represents the position right?

OpenStudy (anonymous):

find the equation of the train given by x(t)=.5t-3.5

myininaya (myininaya):

right x is the position and t is the time

OpenStudy (anonymous):

equation of motion for the train..sorry

myininaya (myininaya):

so we want to know at what time is the position 0?

myininaya (myininaya):

using symbols we say: what does t=? if x=0

OpenStudy (anonymous):

yeah, it tells me to find the position at 4 seconds, the confirm with an equation, then at what instant isthe position zero from your graph and then your equation?

myininaya (myininaya):

well to find when the position is 0 set x=0 and solve for t so we do 0=.5t-3.5 3.5=.5t 2(3.5)=t 7=t

myininaya (myininaya):

we want to know know what x=? whent t=4 so we have x=.5(4)-3.5 x=2-3.5 x=-1.5

OpenStudy (anonymous):

agree with that

myininaya (myininaya):

this second part answers the question: what is the position when the time is 4 seconds

OpenStudy (anonymous):

yeah, i am good on that, its just this, at what instant is the position zero? a) from the graph and b) from your equation.

OpenStudy (anonymous):

i think it would be the same?

myininaya (myininaya):

yes you will get the same answer unless it is hard to read the graph then you might actually get an approximation looking at the graph

OpenStudy (anonymous):

so if the original equation is x(4)=0.5t-3.5 I would just plug in 0 for x (or p) and solve?

myininaya (myininaya):

what is p?

OpenStudy (anonymous):

i am assuming p=position or x axis in this case

myininaya (myininaya):

lol ok lets assume that

myininaya (myininaya):

so if you have x(t)=0.5t-3.5 then x(4)=0.5(4)-3.5

myininaya (myininaya):

which question are you trying to do?

myininaya (myininaya):

the one about t=4 or the one about x=0?

OpenStudy (anonymous):

im gonna scan it in, is that okay?

myininaya (myininaya):

ok

OpenStudy (anonymous):

i almost got it here right now

OpenStudy (anonymous):

it just doesn't seem clear o me what it's asking me to do

myininaya (myininaya):

ok so it looks like you got to e and stop we want to know what time it is when the position is 0 so we want to what t=? when x=0 so we have the equation that relates position and time: x(t)=0.5t-3.5 so we just plug in x=0 0=0.5t-3.5 add 3.5 to both sides 3.5=0.5t multiply 2 on both sides 2(3.5)=t 7=t t=7

OpenStudy (anonymous):

exactly. which is the same thing the graph shows me. so i am assuming they are just wanting me to see the relationship between algebraically and the graph

myininaya (myininaya):

yes

myininaya (myininaya):

they are just trying to get you to see you can look at it both ways

OpenStudy (anonymous):

i can do the next question by plugging x =12m correct?

OpenStudy (anonymous):

12(3.5)=t

OpenStudy (anonymous):

42

myininaya (myininaya):

i didn't get 42

OpenStudy (anonymous):

oops, i forgot the .5t

OpenStudy (anonymous):

it wont cancel because we aren't looking for position =0 position =12

myininaya (myininaya):

we have x(t)=.5t-3.5 we want to know what t=? when x=12 12=.5t-3.5 so solve this equation for t 12+3.5=.5t 15.5=.5t 31=t t=31

myininaya (myininaya):

first step: add 3.5 on both sides second step: multiply 2 on both sides

OpenStudy (anonymous):

i have no idea why i am making this so hard. I know this stuff. Why am I think stuff like this? ( i know you don't know the answer to that lol)

myininaya (myininaya):

lol sometimes you need to take a deep breath and relax for a second and then go back thinking

OpenStudy (anonymous):

im actually in calculus and can't do a simpl y=mx+b problem, it's embarassing

myininaya (myininaya):

i'm teaching cal 2 right now. don't feel bad you can't do algebra. it seems most people still struggle with in cal. i have heard some people said thats where they learned their algebra.

OpenStudy (anonymous):

I just bombed my first CALC exam on limits. and I am in phyiscs. its really important to me that I learn my stuff and get my act together. I want to actually succeed. I am ADD so that screws me over really bad, and I am 28. Embarrassing, cause I really want to learn this stuff well.

myininaya (myininaya):

but you need to work really hard to get caught up i mean i know you probably are doing this already you just need to work really really hard and you will get it

OpenStudy (anonymous):

I don't want to just "get by" i want to actually be able to see the errors in my own problems too.

OpenStudy (anonymous):

oh well, thanks for letting me vent haha. and for all your guidance!

myininaya (myininaya):

so your teacher knows of your learning disorder?

OpenStudy (anonymous):

yes, they all do. However, interestingly enough this hasn't been officially documented. I am going through the testing now, and finish it tomorrow. My Physics Professor caught onto this last year, he has it. My Dr. has always thought I had it, but didn't want to alarm me because I coped so well, until I went back to school that is.

OpenStudy (anonymous):

I notice a major difference on medication, and so does my Prof. just waiting to get this official so I can really start to get some guidance to deal with it better

OpenStudy (anonymous):

this problem, the one you helped me with. i worked on for 1.5 hours. the sad part is, i do know it too!

OpenStudy (anonymous):

my mind wonders, and makes problems harder than they really are. i try to find shortcuts that don't exist, even when I know they don't exist. try to solve problems to fast, have problems thinking slowly through things, ect.

myininaya (myininaya):

well the only advice i have is to work hard (try to relax your mind and don't let your mind go to many places at once) but don't forget to take breaks you can always ask us questions on openstudy.com like you did today

myininaya (myininaya):

good luck to you! you can do it!

OpenStudy (anonymous):

i just found this site too, thanks for your help and advice!

myininaya (myininaya):

np :)